OpenStudy (anonymous):

A bicycle wheel with radius 26" rotates through an arc that measures 80°. What is the length of the arc of the tire that touched the ground?

OpenStudy (anonymous):

Find the circumference

OpenStudy (anonymous):

is this the same ?

OpenStudy (anonymous):

then multiply it by 80/360

OpenStudy (anonymous):

I told you how to do it now you should be able to do it

OpenStudy (anonymous):

104/9?

OpenStudy (anonymous):

or 11.5

OpenStudy (anonymous):

circumference is : \[\Pi2r = \Pi d\]

OpenStudy (anonymous):

\[52 \Pi\]

OpenStudy (anonymous):

Yes, i got that

OpenStudy (anonymous):

yeah but you forgot the pi

OpenStudy (anonymous):

And i did, 52pi*80/360

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

\[52 \Pi \times \frac{ 80 }{ 360 }\]

OpenStudy (anonymous):

see you know how to do it, just put the pi back in

OpenStudy (anonymous):

104/9 * pi

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

\[104\pi/9\]

OpenStudy (anonymous):

The pi times the whole set?

OpenStudy (anonymous):

thats the answer ^^

OpenStudy (anonymous):

the 104 pi / 9

OpenStudy (anonymous):

Okay, thank you AGAIN.